OpenStudy (anonymous):

factor a^2 +16a+28 into binomials. (a-2)(a+14)......(a+2)(a+14)......(a-4)(a-7).....(a+4)(a+7)

OpenStudy (anonymous):

Try yo find two numbers, that sum up to 16 and the result of their product is 28.

OpenStudy (anonymous):

Well, you know that the middle term in the quadratic comes from the sum of the inner and outer terms so if you have (a + k) (a + m) then.. 16 = k + m 28 = k * m

OpenStudy (anonymous):

Solve for k and m.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What did you end up with?